OpenStudy (anonymous):

y-intercept of the line passing through (2,-3) & (-4,1) . I know that the slope is -2/3...

OpenStudy (karatechopper):

ok well do you know what point slope form and slope intercept form is?

OpenStudy (anonymous):

slope intercept y=mx+b

OpenStudy (karatechopper):

and point slope?

OpenStudy (anonymous):

y-y1=m (x-x1)?

OpenStudy (anonymous):

so how do i use this to find the intercept?